OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) Show your workk

OpenStudy (anonymous):

step one: remove parentheses using distributive property \[-2x+6\geq 5-x-3\] step two: combine like terms \[-2x+6\geq 2-x\] step three: add \(2x\) to both sides \[6\geq 2+x\] step four: subtract 2 from both sides \[4\geq x\]

OpenStudy (anonymous):

Im confused about step 2

OpenStudy (anonymous):

Take that back I see what he was talking about. they distributed on the one side and got "5-x-3" . You can add the 5 and -3.

OpenStudy (anonymous):



OpenStudy (anonymous):

That is called combining like terms.

OpenStudy (anonymous):

Does that make sense?